OpenStudy (anonymous):

If 35 percent of b equals 70 percent of d, which of the following expresses d in terms of d?? (A) d=35 percent of b (B) d=50 percent of b (C) d=75 percent of b (D) d=105 percent of b (E) d=200 percent of b

OpenStudy (anonymous):

sorry i messed up the question

OpenStudy (anonymous):

35/100b=70/100d b=70/35d b=2d b=50/100d (B) d=50 percent of b

OpenStudy (saifoo.khan):

You can still edit it. @Studybug123

OpenStudy (anonymous):

If 35 percent of b equals 70 percent of d, which of the following expresses d in terms of b??

OpenStudy (anonymous):

yes I know

OpenStudy (anonymous):

And thank you @sai1234